OpenStudy (anonymous):

if cos feta=-3/5 and pi

OpenStudy (anonymous):

theta

OpenStudy (anonymous):

okok but how do i do this?

OpenStudy (anonymous):

What do you usually work with: degrees, radians?

OpenStudy (anonymous):

radians

OpenStudy (anonymous):

Well this is not from a special angle, and would not be on the trig circle, so you just punch it in your calculator.

OpenStudy (anonymous):

\[\cos \theta =-3/5\]Therefore, the angle theta whose cos is -3/5, stated differently\[\theta =\cos^{-1} (-3/5)\]

OpenStudy (anonymous):

that is the cosine inverse

OpenStudy (anonymous):

Take the inverse of \[\cos \Theta \] and divide it by 2, and you will have \[\Theta/2\]